# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ 3 Results
    The propensity scores are bounded from below, i.e., $\min_{i}\min_{\boldsymbol{x}}\pi_{i}(\boldsymbol{x})\equiv\pi_{\min}>0$.
@@
    The distribution of $\boldsymbol{X}$ is supported on finitely many values, so that $\min_{\boldsymbol{x}}p(\boldsymbol{x})\equiv p_{\min}>0$.
@@
    The CDF $F\equiv F_{Y_{1}}$ is twice continuously differentiable on $[0,1]$.
@@
    The CDF $F_{Y_{1}\mid\boldsymbol{X}_{1}}$ is continuously differentiable on $[0,1]$.
@@
Throughout the paper, $f\equiv F^{\prime}$ denotes the density of $Y$, and similarly $\smash{f_{Y_{1}\mid\boldsymbol{X}_{1}}=F_{Y_{1}\mid\boldsymbol{X}_{1}}^{\prime}}$. The degree parameter $m=m(n)$ is always assumed to be a function of the sample size $n$, and further $m\to\infty$ whenever $n\to\infty$. For real sequences $(U_{n})$ and $(V_{n})$, the notation $U_{n}=\mathcal{O}(V_{n})$ means that $\limsup_{n\to\infty}|U_{n}/V_{n}|\leq C<\infty$, where the nonnegative constant $C$ may depend on $d$, $\pi_{\min}$, $p_{\min}$, $F$, or $F_{Y_{1}\mid\boldsymbol{X}_{1}}$, but no other variable unless explicitly written as a subscript. Typically, the dependence is on the variable $y$, in which case one writes $U_{n}=\mathcal{O}_{y}(V_{n})$. In some places, the Vinogradov notation $U_{n}\ll V_{n}$ is used to indicate $U_{n}=\mathcal{O}(V_{n})$ more concisely. Similarly, the notation $U_{n}=\mathrm{o}(V_{n})$ means that $\lim_{n\to\infty}|U_{n}/V_{n}|=0$. Subscripts indicate which parameters the convergence rate can depend on. More generally, if $U_{n},V_{n}$ are random variables, we write $U_{n}=\mathcal{O}_{L^{2}}(V_{n})$ if $\limsup_{n\to\infty}\mathsf{E}[|U_{n}/V_{n}|^{2}]\leq C<\infty$ and $U_{n}=\mathrm{o}_{L^{2}}(V_{n})$ if $\lim_{n\to\infty}\mathsf{E}[|U_{n}/V_{n}|^{2}]=0$. If the rate depends on a variable such as $y$, we add a subscript.
@@ 3.1 Asymptotic properties of $\smash{\widetilde{F}_{n,m}}$
The following proposition establishes the asymptotics of the pointwise bias. Since the underlying pseudo-empirical estimator $\widetilde{F}_{n}$ is unbiased for $F$, the bias of the smoothed estimator $\smash{\widetilde{F}_{n,m}}$ is entirely attributable to the Bernstein smoothing process, confirming the standard bias rate of $\mathcal{O}(m^{-1})$.
@@
Suppose that Assumption (A3) holds. Then, uniformly for $y\in[0,1]$, we have, as $n\to\infty$,
$\mathsf{Bias}(\widetilde{F}_{n,m})=\mathsf{E}[\widetilde{F}_{n,m}(y)]-F(y)=m^{-1}B(y)+\mathrm{o}(m^{-1}),$
@@
Next, we examine the pointwise variance of the pseudo estimator. The expansion reveals two main components: the leading term $n^{-1}\sigma^{2}(y)$, corresponding to the variance of the unsmoothed IPW empirical CDF, and a negative correction term of order $\mathcal{O}(n^{-1}m^{-1/2})$. This correction term explicitly quantifies the variance reduction achieved by the Bernstein smoothing.
@@
$\sigma^{2}(y)=\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}\right]-\{F(y)\}^{2},\qquad V(y)=\sqrt{\frac{y(1-y)}{\pi}}\mathsf{E}\left[\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}\right].$ (3.2)
@@
$=n^{-1}\sigma^{2}(y)-n^{-1}m^{-1/2}V(y)+m^{-2}B^{2}(y)+\mathrm{o}_{y}(n^{-1}m^{-1/2})+\mathrm{o}(m^{-2}).$
@@
To obtain a global measure of accuracy, we integrate the MSE over the support $[0,1]$. This yields the MISE and allows for the determination of a globally optimal degree $m_{\mathrm{opt}}$. This optimal degree achieves a convergence rate of $n^{-4/3}$ for the second-order term, demonstrating a clear improvement over the unsmoothed estimator $\smash{\widetilde{F}_{n}}$.
@@
Next, we establish the asymptotic normality of the pseudo estimator, which is essential for constructing confidence intervals. The proof is a straightforward verification of the Lindeberg condition for double arrays.
@@
$n^{1/2}\{\widetilde{F}_{n,m}(y)-F(y)\}\stackrel{{\scriptstyle\mathrm{law}}}{{\longrightarrow}}\begin{cases}\mathcal{N}(0,\sigma^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to 0,\\ \mathcal{N}(\lambda B(y),\sigma^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to\lambda.\end{cases}$
@@ 3.2 Asymptotic properties of $\smash{\widehat{F}_{n,m}}$
We now turn to the feasible estimator $\smash{\widehat{F}_{n,m}}$, which uses estimated propensity scores. We analyze its properties by characterizing its relationship with the pseudo estimator $\smash{\widetilde{F}_{n,m}}$. We first show that the process of estimating the propensity scores introduces only a negligible amount of additional bias of order $\mathcal{O}(n^{-1})$, which can be positive or negative.
@@
Suppose that Assumptions (A1)(A3) hold. Then, uniformly for $y\in[0,1]$, we have, as $n\to\infty$,
@@
A key finding is that the feasible estimator exhibits a notable variance reduction compared to the pseudo estimator. Proposition 7 reveals that $\smash{\widehat{F}_{n,m}}$ has a smaller asymptotic variance than $\smash{\widetilde{F}_{n,m}}$ by an explicit nonnegative correction term $n^{-1}C(y)$. This highlights a beneficial interaction between smoothing and nonparametric propensity estimation in this context. Let us remark that Dubnicka, (2009, Eq. (8)) obtained a similar variance reduction in the KDE setting.
@@
$C(y)$ $=\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right],$ (3.3)
@@
$=\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)(1-F_{Y_{1}\mid\boldsymbol{X}_{1}}(y))}{\pi_{1}(\boldsymbol{X}_{1})}\right]+\mathsf{E}\left[\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]-\{F(y)\}^{2}.$
@@
$\mathsf{Var}(\widehat{F}_{n,m}(y))$ $=\mathsf{Var}(\widetilde{F}_{n,m}(y))-n^{-1}C(y)+\mathcal{O}(n^{-1}m^{-1})$
@@
$=\mathrm{MSE}[\widetilde{F}_{n,m}(y)]-n^{-1}C(y)+\mathcal{O}(n^{-1}m^{-1})+\mathcal{O}(n^{-2})$
$=n^{-1}\nu^{2}(y)-n^{-1}m^{-1/2}V(y)+m^{-2}B^{2}(y)+\mathrm{o}_{y}(n^{-1}m^{-1/2})+\mathrm{o}(m^{-2})+\mathcal{O}(n^{-2}).$
@@
Finally, we establish the asymptotic normality of the feasible estimator. The conditions on the polynomial degree remain the same as for the pseudo estimator, but the asymptotic variance $\nu^{2}(y)$ is smaller than $\sigma^{2}(y)$, reflecting the efficiency gain from estimating the propensity scores.
@@
$n^{1/2}\{\widehat{F}_{n,m}(y)-F(y)\}\stackrel{{\scriptstyle\mathrm{law}}}{{\longrightarrow}}\begin{cases}\mathcal{N}(0,\nu^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to 0,\\ \mathcal{N}(\lambda B(y),\nu^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to\lambda.\end{cases}$
@@ 4 Monte Carlo simulations
In this section, we conduct a modest Monte Carlo study to evaluate three families of estimators for the CDF under MAR: the unsmoothed IPW empirical CDF, its Bernstein-smoothed version, and the integrated IPW Gaussian KDE of Dubnicka, (2009) (abbreviated I-IPW KDE). Each family is examined in two regimes: a pseudo version using known propensities and a feasible version using propensities estimated nonparametrically, as described in Section 2. All simulations were implemented in R; see A for the GitHub link to the code.
@@ 4.1 Data generating process and setup
We consider a setting with a univariate continuous response $Y$ and a univariate discrete auxiliary covariate $X$ (d=1d{=}1). This entails no loss of generality for discrete covariates, given that any finite-dimensional discrete $\boldsymbol{X}$ can be recoded into a single factor by labeling the cells of the Cartesian product of its marginal categories. The data generating process used in the code is:
  • $X_{i}\sim\mathrm{Bernoulli}(p_{X})$ with $p_{X}=0.5$;
  • $Y_{i}\sim\mathrm{Beta}(\alpha,\beta)$ with $(\alpha,\beta)=(2,5)$;
  • Missingness is MAR with propensity $\pi(X_{i})$ given by
    $\pi(x)=\begin{cases}0.6,&x=0,\\ 0.9,&x=1,\end{cases}\qquad\text{so that }~\mathsf{E}[\pi(X_{i})]=0.75\text{ and }\pi_{\min}=0.6.$
@@
  3. (iii)
    the pseudo I-IPW KDE, obtained by integrating the IPW KDE of Dubnicka, (2009) constructed with $W_{i}$;
@@
    the Bernstein-smoothed feasible estimator $\widehat{F}_{n,m}=\mathcal{B}_{m}(\widehat{F}_{n})$ (smoothing $\widehat{F}_{n}$);
@@
    the feasible I-IPW KDE, obtained by integrating the IPW KDE built with $\widehat{W}_{i}$.
@@
We use sample sizes $n\in\{25,50,100,200,400,800,1600,3200,6400\}$ and $N=100$ Monte Carlo replications per $n$. For each replication, we compute the integrated squared error (ISE),
@@
numerically via a uniform Riemann sum over $G=2^{10}=512$ equally spaced grid points on $[0,1]$. We summarize the distribution of the ISE across replications by its mean, median, interquartile range, and variance. Replications are parallelized over the available CPU cores.
4.2 Bandwidth selection via optimized LSCV

### v2

@@ 3 Results
    The propensity scores are bounded from below, i.e., $\pi_{\min}\vcentcolon=\min_{\boldsymbol{x}}\pi(\boldsymbol{x})>0$.
@@
    The distribution of $\boldsymbol{X}$ is supported on finitely many values, so that $p_{\min}\vcentcolon=\min_{\boldsymbol{x}}p(\boldsymbol{x})>0$.
@@
    The CDF $F$ (equivalently, $F_{Y_{1}}$) is three times continuously differentiable on $(0,1)$.
@@
    The CDF $F_{Y_{1}\mid\boldsymbol{X}_{1}}$ is twice continuously differentiable on $(0,1)$.
@@
Throughout the paper, $f$ (equivalently, $f_{Y_{1}}$) denotes the density of $Y_{1}$, and similarly $\smash{f_{Y_{1}\mid\boldsymbol{X}_{1}}}$ denotes the conditional density of $Y_{1}$ given $\boldsymbol{X}_{1}$. Since $Y_{1}$ is a continuous random variable, note that $f=F^{\prime}$ and $\smash{f_{Y_{1}\mid\boldsymbol{X}_{1}}=F_{Y_{1}\mid\boldsymbol{X}_{1}}^{\prime}}$. The degree parameter $m=m(n)$ is always assumed to be a function of the sample size $n$, and further $m\to\infty$ whenever $n\to\infty$. For real sequences $(U_{n})$ and $(V_{n})$, the notation $U_{n}=\mathcal{O}(V_{n})$ means that $\limsup_{n\to\infty}|U_{n}/V_{n}|\leq C<\infty$, where the nonnegative constant $C$ may depend on $d$, $\pi_{\min}$, $p_{\min}$, $F$, or $F_{Y_{1}\mid\boldsymbol{X}_{1}}$, but no other variable unless explicitly written as a subscript. Typically, the dependence is pointwise on the variable $y$, in which case one writes $U_{n}=\mathcal{O}_{y}(V_{n})$. More specifically, $U_{n}=\mathcal{O}_{y}(V_{n})$ means that for any fixed $y$, there exists $C_{y}>0$ such that $\limsup_{n\to\infty}|U_{n}/V_{n}|\leq C_{y}<\infty$. In some places, the Vinogradov notation $U_{n}\ll V_{n}$ is used to indicate $U_{n}=\mathcal{O}(V_{n})$ more concisely. Similarly, the notation $U_{n}=\mathrm{o}(V_{n})$ means that $\lim_{n\to\infty}|U_{n}/V_{n}|=0$. Subscripts indicate which parameters the convergence rate can depend on. More generally, if $U_{n},V_{n}$ are random variables, we write $U_{n}=\mathcal{O}_{L^{2}}(V_{n})$ if $\limsup_{n\to\infty}\mathsf{E}[|U_{n}/V_{n}|^{2}]\leq C<\infty$ and $U_{n}=\mathrm{o}_{L^{2}}(V_{n})$ if $\lim_{n\to\infty}\mathsf{E}[|U_{n}/V_{n}|^{2}]=0$. If the rate depends on a variable such as $y$, we add a subscript.
@@ 3.1 Asymptotic properties of $\smash{\widetilde{F}_{n,m}}$
The following proposition establishes the asymptotics of the pointwise bias. Since the underlying pseudo-empirical estimator $\widetilde{F}_{n}$ is unbiased for $F$, the bias of the smoothed estimator $\smash{\widetilde{F}_{n,m}}$ is entirely attributable to the Bernstein smoothing process, confirming the standard bias rate of order $m^{-1}$.
@@
Suppose that Assumptions (A1) and (A3) hold. Then, for any $y\in(0,1)$, we have, as $n\to\infty$,
$\mathsf{Bias}(\widetilde{F}_{n,m}(y))=\mathsf{E}[\widetilde{F}_{n,m}(y)]-F(y)=m^{-1}B(y)+\mathrm{o}_{y}(m^{-1}),$
@@
Next, we examine the pointwise variance of the pseudo estimator. The expansion reveals two main components: the leading term $n^{-1}\sigma^{2}(y)$, corresponding to the variance of the unsmoothed IPW empirical CDF, and a negative correction term of order $n^{-1}m^{-1/2}$. This correction term explicitly quantifies the variance reduction achieved by the Bernstein smoothing.
@@
$\sigma^{2}(y)=\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}\right]-\{F(y)\}^{2},\qquad V(y)=\sqrt{\frac{y(1-y)}{\pi}}\mathsf{E}\left[\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}\right].$ (3.2)
@@
$=n^{-1}\sigma^{2}(y)-n^{-1}m^{-1/2}V(y)+m^{-2}B^{2}(y)+\mathrm{o}_{y}(n^{-1}m^{-1/2})+\mathrm{o}_{y}(m^{-2}).$
@@
To obtain a global measure of accuracy, we integrate the MSE over $(0,1)$. This yields the MISE and allows for the determination of a globally optimal degree $m_{\mathrm{opt}}$. This optimal degree achieves a convergence rate of $n^{-4/3}$ for the second-order term, demonstrating a clear improvement over the unsmoothed estimator $\smash{\widetilde{F}_{n}}$.
@@
Next, we establish the asymptotic normality of the pseudo estimator, which is essential for constructing pointwise confidence intervals. The proof is a straightforward verification of the Lindeberg condition for double arrays. Developing uniform confidence bands for the entire curve, however, falls outside the scope of this paper and would typically require bootstrap or multiplier methods.
@@
$n^{1/2}\{\widetilde{F}_{n,m}(y)-F(y)\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\begin{cases}\mathcal{N}(0,\sigma^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to 0,\\ \mathcal{N}(\lambda B(y),\sigma^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to\lambda,\end{cases}$
where $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes the convergence in distribution (pointwise in $y$).
@@ 3.2 Asymptotic properties of $\smash{\widehat{F}_{n,m}}$
We now turn to the feasible estimator $\smash{\widehat{F}_{n,m}}$, which uses estimated propensity scores. Throughout this subsection, we work under the discrete finite-support assumption on $\boldsymbol{X}$ from Section 2. We analyze its properties by characterizing its relationship with the pseudo estimator $\smash{\widetilde{F}_{n,m}}$. We first show that the process of estimating the propensity scores introduces only a negligible amount of additional bias of order $n^{-1}$, which can be positive or negative.
@@
A key finding is that the feasible estimator exhibits a notable variance reduction compared to the pseudo estimator. Proposition 7 reveals that $\smash{\widehat{F}_{n,m}}$ has a smaller asymptotic variance than $\smash{\widetilde{F}_{n,m}}$ by an explicit nonnegative correction term $n^{-1}C(y)$. This highlights a beneficial interaction between smoothing and nonparametric propensity estimation in this context. Let us remark that Dubnicka (2009, Eq. (8)) obtained a similar variance reduction in the KDE setting.
@@
$C(y)$ $=\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right],$ (3.3)
@@
$=\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)(1-F_{Y_{1}\mid\boldsymbol{X}_{1}}(y))}{\pi(\boldsymbol{X}_{1})}\right]+\mathsf{E}\left[\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]-\{F(y)\}^{2}.$
@@
$\mathsf{Var}(\widehat{F}_{n,m}(y))$ $=\mathsf{Var}(\widetilde{F}_{n,m}(y))-n^{-1}C(y)+\mathcal{O}_{y}(n^{-1}m^{-1})$
@@
Remark 2.
The variance reduction in Proposition 7 does not rely on the auxiliary variable $\boldsymbol{X}$ being discrete with finite support. As mentioned in Section 2, the restriction to the discrete case was adopted mainly to avoid the tedious task of treating the discrete and continuous covariate settings in parallel. The same phenomenon was proved by Dubnicka (2009) for the IPW KDE with continuous covariates. In the present CDF setting with continuous covariates, the same asymptotic results hold if the propensity scores are estimated by a Nadaraya–Watson estimator.
@@
$=\mathrm{MSE}[\widetilde{F}_{n,m}(y)]-n^{-1}C(y)+\mathcal{O}_{y}(n^{-1}m^{-1})+\mathcal{O}(n^{-2})$
$=n^{-1}\nu^{2}(y)-n^{-1}m^{-1/2}V(y)+m^{-2}B^{2}(y)+\mathrm{o}_{y}(n^{-1}m^{-1/2})+\mathrm{o}_{y}(m^{-2})+\mathcal{O}(n^{-2}).$
@@
Finally, we establish the asymptotic normality of the feasible estimator. The conditions on the polynomial degree remain the same as for the pseudo estimator, but the asymptotic variance $\nu^{2}(y)$ is smaller than $\sigma^{2}(y)$, reflecting the efficiency gain from estimating the propensity scores. As is the case with the pseudo estimator, deriving functional limit results for this feasible process to perform uniform inference would necessitate further techniques, such as the bootstrap.
@@
$n^{1/2}\{\widehat{F}_{n,m}(y)-F(y)\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\begin{cases}\mathcal{N}(0,\nu^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to 0,\\ \mathcal{N}(\lambda B(y),\nu^{2}(y)),&\mbox{if }n^{1/2}m^{-1}\to\lambda,\end{cases}$
where again the convergence in distribution is pointwise in $y$.
@@ 4 Monte Carlo simulations
In this section, we conduct a modest Monte Carlo study to evaluate three families of estimators for the CDF under MAR: the unsmoothed IPW empirical CDF, its Bernstein-smoothed version, and the integrated IPW Gaussian KDE of Dubnicka (2009) (abbreviated I-IPW KDE). Each family is examined in two regimes: a pseudo version using known propensities and a feasible version using propensities estimated nonparametrically, as described in Section 2. All simulations were implemented in R; see Appendix A for the GitHub link to the code.
@@ 4.1 Data generating process and setup
We consider a setting with a univariate continuous response $Y$ and a univariate discrete auxiliary covariate $X$ ($d=1$). This entails no loss of generality for discrete covariates, given that any finite-dimensional discrete $\boldsymbol{X}$ can be recoded into a single factor by labeling the cells of the Cartesian product of its marginal categories. The data-generating process is:
  • $Y_{i}\sim\mathrm{Beta}(\alpha,\beta)$ with $(\alpha,\beta)=(0.9,0.9)$;
  • letting $F_{\mathrm{Beta}(0.9,0.9)}$ denote the $\mathrm{Beta}(0.9,0.9)$ CDF and $\Phi$ the standard normal CDF, define
    $U_{i}=F_{\mathrm{Beta}(0.9,0.9)}(Y_{i}),\qquad T_{i}=\Phi^{-1}(U_{i}),$
    so that $T_{i}\sim\mathcal{N}(0,1)$. Next, generate
    $X_{i}^{\star}=\rho\,T_{i}+\sqrt{1-\rho^{2}}\,Z_{i},\qquad Z_{i}\sim\mathcal{N}(0,1),$
    with $Z_{i}$ independent of $Y_{i}$, and fix $\rho=0.6$. Finally, define the discrete auxiliary covariate by
    $X_{i}=\mathds{1}_{\{X_{i}^{\star}>0\}}.$
    Then $X_{i}\sim\mathrm{Bernoulli}(p_{X})$ with $p_{X}=0.5$. In particular, $X_{i}$ and $Y_{i}$ are dependent through the latent Gaussian construction.
  • missingness is MAR with logistic propensity
    $\pi(x)=\mathsf{P}(\delta_{i}=1\mid X_{i}=x)=[1+\exp\{-(\beta_{0}+\beta_{1}x)\}]^{-1},\qquad x\in\{0,1\},$
    where $(\beta_{0},\beta_{1})$ are chosen so that
    $\pi(0)=0.6,\qquad\pi(1)=0.9.$
    Equivalently,
    $\beta_{0}=\log(0.6/0.4),\qquad\beta_{1}=\log(0.9/0.1)-\log(0.6/0.4),$
    so that $\mathsf{E}[\pi(X_{i})]=0.75$ and $\pi_{\min}=0.6$.
@@
    the pseudo I-IPW KDE, obtained by integrating the IPW KDE of Dubnicka (2009) constructed with $W_{i}$;
  3. (iii)
@@
    the feasible I-IPW KDE, obtained by integrating the IPW KDE built with $\widehat{W}_{i}$;
@@
    the Bernstein-smoothed feasible estimator $\widehat{F}_{n,m}=\mathcal{B}_{m}(\widehat{F}_{n})$ (smoothing $\widehat{F}_{n}$).
@@
We conduct two Monte Carlo experiments, each with $1000$ replications. The first (Section 4.3) examines the impact of the sample size under the benchmark MAR mechanism defined above, with
$n\in\{25,50,100,200,400,800,1600,3200,6400\}.$
The second (Section 4.4) examines the impact of the missing rate at the fixed sample size $n=400$. In that experiment, we retain the same latent Gaussian dependence between $X$ and $Y$, and we vary only the intercept of the logistic MAR mechanism, keeping the slope coefficient $\beta_{1}$ fixed at its benchmark value. More precisely, for each target missing rate $r\in\{5\%,10\%,15\%,20\%,25\%,30\%,35\%,40\%\}$, we define
$\pi_{r}(x)=\{1+\exp[-(\beta_{0}(r)+\beta_{1}x)]\}^{-1},\qquad x\in\{0,1\},$
where $\beta_{0}(r)$ is chosen so that
$1-\mathsf{E}[\pi_{r}(X_{i})]=r.$
Performance measures
For each replication, we compute the integrated squared error,
@@
and the boundary ISE,
$\mathrm{BISE}(\widehat{F})=\frac{1}{2\delta}\int_{[0,\delta]\cup[1-\delta,1]}\{\widehat{F}(y)-F(y)\}^{2}\,\mathrm{d}y,\qquad\delta=n^{-2/3}.$
For each estimator, we compute the integrals numerically using the cubature routine adaptIntegrate. We summarize the distribution of the ISEs and BISEs across replications by their mean and standard deviation.
4.2 Smoothing-parameter selection via optimized LSCV